OpenStudy (anonymous):

Solve: (1/6x^2 ) = ( 1/3x^2) - (1/x)

OpenStudy (anonymous):

x=1/6

OpenStudy (anonymous):

i can show work if needed

OpenStudy (anonymous):

multiply through by the lcm 6x^2: 1 = 2 - 6x solve for x

OpenStudy (anonymous):

@Muqn yes could you explain.

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

its hard for me to show my work the way i did it on paper mabe jowywhite can better

OpenStudy (anonymous):

joeywhite

OpenStudy (anonymous):

ok thats fine.

OpenStudy (anonymous):

ok sorry just cant do it on here lol

OpenStudy (anonymous):

a fractional equation can always be solved by multiplying by lcm of the denominators 6x^2 , 3x^2 and x the lcm (lowest common multiple) is the lowest value into which these 3 terms can divide lcm of 3 and 6 is 6 lcm of x^2, x is x^2 so the lcm of all 3 is 6 x^2 now do the multiplication: